Consumption =115+0.6y
Investment = 550

4.3 Calculate the equilibrium level of income
1
Expert's answer
2020-09-25T09:19:48-0400

For equilibrium level of income to occur, aggregate supply (AS) = aggregate demand (AD).

let "Y" be income in the economy.

"Y" = "C" + "I" where "C" is the consumption, and "I" is the investment

"Y = 115 + 0.6Y + 550"

"Y - 0.6Y = 115 + 550"

"0.4Y = 665"

"Y = \frac{665}{0.4} = 1662.5"